countryboy3642:

At 8:45 p.m., a jet is located 64 mi due east of a city. A second jet is located 57 mi due north of the city. To the nearest tenth of a mile, what is the distance between the two jets? Enter your answer as a decimal in the box. mi

eatsalad:

a^2+b^2=c^2 Where a is 57 and b is 64 57^2+64^2=c^2 Solve. and get 85.70